# zkPHIRE: A Programmable Accelerator for ZKPs over HIgh-degRee, Expressive Gates

### v1

@@ Abstract
Zero-Knowledge Proofs (ZKPs) have emerged as powerful tools for secure and privacy-preserving computation. ZKPs enable one party to convince another of a statement’s validity without revealing anything else. This capability has profound implications in many domains, including: machine learning, blockchain, image authentication, and electronic voting. Despite their potential, ZKPs have seen limited deployment because of their exceptionally high computational overhead, which manifests primarily during proof generation. To mitigate these overheads, a (growing) body of researchers has proposed hardware accelerators and GPU implementations for kernels and complete protocols. Prior art spans a wide variety of ZKP schemes that vary significantly in computational overhead, proof size, verifier cost, protocol setup, and trust. The latest, and widely used ZKP protocols are intentionally designed to balance these trade-offs. A particular challenge in modern ZKP systems is supporting complex, high-degree gates using the SumCheck protocol. We address this challenge with a novel programmable accelerator that efficiently handles arbitrary custom gates via SumCheck. Our accelerator achieves upwards of $1000\times$ geomean speedup over CPU-based SumChecks across a range of gate types. We integrate this unit into a full-system accelerator, zkPHIRE, which achieves $1486\times$ geomean speedup over CPU and $11.87\times$ speedup over the state-of-the-art at iso-area. zkPHIRE is the first accelerator to scale to problem sizes of $2^{30}$ nominal constraints while maintaining small proof sizes and programmability.
**footnotetext: These authors contributed equally to this work
@@ I Introduction
Zero-Knowledge Proofs (ZKPs) are cryptographic protocols that allow a prover to convince one or more verifiers that a claim is correct without revealing any information beyond the claim’s validity. For example, a prover can demonstrate possession of a valid authentication key without disclosing the key. ZKPs have many compelling applications, including privacy-preserving authentication, trustworthy AI, blockchain scalability, and recently, private age verification [53]. A major limitation of state-of-the-art ZKP protocols is their high computational cost–generating a single proof on a CPU can take minutes to hours. As a result, there is growing interest in accelerating ZKPs, both with programmable solutions (e.g., GPUs [36, 24]) and ASICs [66, 13, 67, 49, 12].
@@
ZKP protocols continue to evolve and improve. Many have been proposed, offering different trade-offs in runtime, proof size, verifier time, and setup (universal vs. application-specific), and security guarantees. A major family of ZKP protocols relies on the SumCheck protocol[35]. SumCheck-based protocols have $O(N)$ prover time, as compared to NTT-based protocols (a popular alternative[20, 17]), which require $O(N\log(N))$ prover operations. The SumCheck protocol is highly flexible and is at the core of many classic and modern ZKP protocols [19, 50, 56, 9, 61, 6, 64], which have relatively fast prover time. While SumChecks were part of previously accelerated systems [55, 49, 12], they studied specific instantiations of the SumCheck protocol rather than the protocol family more generally. Another major distinguishing factor between protocols is the type of polynomial commitment scheme[25] that is used. Some of these protocols, such as Spartan [50]+Orion [61] use faster Merkle tree commitments, at the cost of large proofs (many MBs), while HyperPlonk [9] uses a slower commitment scheme that produces significantly smaller proofs (a few kBs).
@@
ZKP protocols typically encode computations as arithmetic circuits (ACs) – directed acyclic graphs where each node (or gate) enforces a local algebraic constraint, such as addition ($f=a+b$) or multiplication ($f=a\cdot b$) over a finite field. We generally denote the number of gates by $N$. These circuits are then arithmetized by translating gates into polynomial constraints, which are later verified using polynomial checks for correctness. The SumCheck protocol is used to reduce checking $N$ gates to checking a single, randomized gate formula. A key feature of newer SumCheck-based protocols is the use of custom, high degree gates that encapsulate more computation per gate by using higher-degree polynomial constraints (e.g., $f=a\cdot b^{5}+c^{2}$). This reduces the overall number of gates needed to represent a computation, resulting in large efficient gates. The move to custom gates is exemplified in the Halo2 arithmetization language[63]. Halo2 enables a circuit designer to specify the types of gates they want to use in each circuit.
@@
However, this level of customizability introduces significant hardware challenges. Unlike NTTs, which have fixed, regular dataflows, SumCheck computations vary with each gate’s structure and polynomial degree. High-degree polynomials require more modular multiplications and deeper reductions, leading to increased bandwidth pressure, varying memory access patterns, and more complex scheduling. Designing efficient SumCheck hardware thus involves a tradeoff: either build fixed-function units for specific gate shapes, or develop a general, programmable architecture that can dynamically adapt to diverse polynomials without significantly compromising performance.
@@
zkSpeed [12] exemplifies the first approach, using dedicated hardware for Vanilla gates (additions and multiplications) within the HyperPlonk protocol. While this provides good performance for a fixed gate set, it limits flexibility: zkSpeed’s SumCheck unit cannot easily handle complex custom gates like those seen in the Halo2 [63] arithmetization, reducing its applicability across general ZKP workloads. Furthermore, zkSpeed’s reliance on large on-chip scratchpads limits scalability, as its memory footprint grows with the number of gates, becoming impractical beyond $2^{24}$ gates, a scale increasingly common in emerging ZKP applications.
@@
Generalizable SumCheck cores are needed beyond ZKPs, too. In verifiable computing, SumCheck serves as the core verification mechanism in protocols that provide integrity guarantees without requiring zero knowledge. However, previous work in this area, including verifiable ASICs [55], has also relied on custom SumCheck units with fixed functions, again restricting applicability and reuse.
@@
In this work, we adopt the second approach with zkPHIRE, introducing a programmable SumCheck accelerator that supports arbitrary polynomial structures and gate types. zkPHIRE addresses the irregularity and data reuse challenges intrinsic to SumCheck with fused compute pipelines, tree-based interconnects for efficient reductions, and flexible scheduling to handle diverse polynomial structures and degrees. This flexibility opens the door to a wide range of applications across ZKPs and verifiable computing. We evaluate our programmable SumCheck unit on a diverse array of gate types, and then embed it into a full-system accelerator, zkPHIRE. zkPHIRE improves upon zkSpeed with more efficient memory organization and module abstraction that allows hardware to be reused across protocol steps. We conduct a protocol-level evaluation using the HyperPlonk with custom gates as a representative case study. Our key contributions are:
@@
  • The first hardware evaluation of high-degree gates that offer significant performance and scalability benefits
@@ II-B Multi-Scalar Multiplication
Multi-Scalar multiplication (MSM) is a key kernel used in polynomial commitments and relies on elliptic curve cryptography. MSMs involve computing a linear combination (i.e., dot product) of scalars and corresponding (2D or 3D) points on an elliptic curve. Given scalars $k_{1},k_{2},\ldots,k_{n}$ and corresponding elliptic curve points $P_{1},P_{2},\ldots,P_{n}$, the goal is to efficiently compute $s=\sum_{i}k_{i}P_{i}$. MSMs are computationally intensive because each term $k_{i}P_{i}$ involves a point-scalar multiplication – a costly elliptic curve operation. MSMs constitute a significant portion of the total runtime in many zkSNARKs, including Groth16 and HyperPlonk. As a result, prior work has focused on accelerating MSMs using both application-specific and programmable platforms [13, 66, 34, 36, 24, 30, 67]. Rather than computing each scalar multiplication directly, optimized approaches such as Pippenger’s algorithm [45] are commonly employed. Pippenger’s algorithm improves efficiency by restructuring the computation to rely more heavily on faster operations like point addition and point doubling.
@@ II-C SumCheck
SumCheck [35] is a round-based Interactive Proof (IP) protocol where a prover, $\mathcal{P}$, attempts to convince a verifier, $\mathcal{V}$, that the sum of a multivariate polynomial, $f(X_{1},\ldots,X_{\mu})$, over the boolean hypercube (all boolean values of $X_{i}\ldots X_{\mu}$) equals a claimed value, $C$. The protocol proceeds in $\mu$ rounds, one for each variable of the polynomial. In each round $i$, $\mathcal{P}$ sends $\mathcal{V}$ a univariate polynomial:
@@
where $r_{1},\ldots,r_{i-1}$ are $\mathcal{V}$’s random challenges from prior rounds. $\mathcal{V}$ checks that the sum of $s_{i}(0)+s_{i}(1)$ equals the value claimed in the previous round (or the initial claim $C$ for the first round). $\mathcal{V}$ then chooses a random field element $r_{i}$ and sends it to $\mathcal{P}$, fixing the variable $X_{i}=r_{i}$ for the next round. After $\mu$ rounds, $\mathcal{V}$ evaluates $f$ at the final point $(r_{1},\ldots,r_{\mu})$. If all checks pass, $\mathcal{V}$ accepts.
ZKPs typically execute the SumCheck protocol over multilinear extensions (MLEs) – polynomials with degree at most 1 in each variable. MLEs can represent multivariate functions compactly and are well-suited for hardware because they can be stored and accessed as flat lookup tables indexed by binary-encoded inputs. Throughout this paper, we will use the terms “MLEs”, “MLE tables,” and “polynomials” interchangeably.
@@ II-C1 Polynomials and Gates
Polynomials are important in ZKP protocols because they encode and help enforce the correctness of the computation being proven. A popular encoding scheme is Plonk [17]. Plonk’s arithmetization scheme converts the underlying computation into an arithmetic circuit consisting of gates. Plonk circuits support additions, multiplications, conditionality, and equality checks. They can be represented by the composite polynomial
$f=q_{L}w_{1}+q_{R}w_{2}+q_{M}w_{1}w_{2}-q_{O}w_{3}+q_{C}$ (1)
consisting of individual polynomials $q_{i}$ and $w_{i}$. These are known as selectors (or enables) and witnesses, respectively. Each of these polynomials encodes a function of a multivariate input vector $X=[X_{1},X_{2},\ldots X_{\mu}]$, which encodes the gate index in binary. For example, if $\mu=3$, there are 8 gates, and gate index 3 is represented by evaluating $f(1,1,0)$. For a multiply operation, we would set $q_{M}(1,1,0)=1$, $q_{O}(1,1,0)=1$ and the other enable polynomials to 0. Then, gate 3 would execute correctly if and only if $f(1,1,0)=0$, that is, $w_{1}(1,1,0)\cdot w_{2}(1,1,0)-w_{3}(1,1,0)=0$. Representing circuits as polynomials naturally suits SumCheck-based provers where we can check (with additional constraints) that each gate evaluates to 0 and that $\sum_{X}f(X)=0$.
@@ II-C2 High-degree, Expressive Gates
The challenge of using Plonk gates is that all computations must be mapped to only what is supported by the Plonk polynomial, similar to mapping program code to a RISC ISA. In some cases, this results in circuits that require millions ($2^{20}$) of gates, and in turn, polynomials with millions of evaluations.
[…]
@@ II-C3 SumCheck on Multilinear Polynomial Products
SumCheck over a single polynomial, such as $\sum f(X_{1},X_{2},X_{3})$ (with $\mu=3$) is relatively straightforward. Using the equation for $s_{1}(X_{1})$ from Section II-C, and viewing $f(X)$ in its MLE table form, in round 1, $s_{1}(0)$ is the sum of all even indices (where $X_{1}=0$), and $s_{1}(1)$ is the sum of all odd indices ($X_{1}=1$). This gives us two evaluations of the univariate, linear polynomial $s_{1}(X_{1})$ to be sent to the verifier. We then perform an MLE Update by fixing $X_{1}$ to random challenge $r_{1}$. At fixed $X_{2}$, $X_{3}$, entries $(0,X_{2},X_{3})$ and $(1,X_{2},X_{3})$ represent a line over $X_{1}$, so fixing $X_{1}=r_{1}$ means extrapolating the line to $r_{1}$: $f(r_{1},X_{2},X_{3})=f(0,X_{2},X_{3})\cdot(1-r_{1})+f(1,X_{2},X_{3})\cdot r_{1}$. This operation halves the size of each MLE table.
SumCheck over polynomials using Plonkish encodings (like Equation 1) is more complicated because there are multiple terms with products of polynomials. Figure 1 shows the dataflow for a single polynomial product term. For a $d$-degree polynomial, we require $d+1$ evaluations for $s_{i}(X_{i})$ after each round of SumCheck. This is done by evaluating all terms at $d+1$ points. Each pair of entries in the MLE table are actually evaluations at $X_{i}=0$ and $1$; these values are then extended to $X_{i}=2,\ldots,d$, using the same formula as MLE Update. Then, the extensions for all polynomials in a term are multiplied together. These products are summed (accumulated) down the MLE table, yielding $d+1$ values for a given term. When there are more than one term, this process will be repeated across terms with the $d+1$ evaluations summed across terms to form the final polynomial $s_{i}(X_{i})$ for round $i$.

### v2

@@ Abstract
Zero-Knowledge Proofs (ZKPs) have emerged as a powerful tool for secure and privacy-preserving computation. ZKPs enable one party to convince another of a statement’s validity without revealing anything else. This capability has profound implications in many domains, including machine learning, blockchain, image authentication, and electronic voting. Despite their potential, ZKPs have seen limited deployment because of their exceptionally high computational overhead, which manifests primarily during proof generation. To mitigate these overheads, a (growing) body of researchers has proposed hardware accelerators and GPU implementations of both kernels and complete protocols. Prior art spans a wide variety of ZKP schemes that vary significantly in computational overhead, proof size, verifier cost, protocol setup, and trust. The latest and widely used ZKP protocols are intentionally designed to balance these trade-offs. One particular challenge in modern ZKP systems is supporting complex, high-degree gates using the SumCheck protocol. We address this challenge with a novel programmable accelerator to efficiently handle arbitrary custom gates via SumCheck. Our accelerator achieves upwards of $1000\times$ geomean speedup over CPU-based SumChecks across a range of gate types. We include this unit in zkPHIRE, a programmable, full-system accelerator that accelerates the HyperPlonk protocol. zkPHIRE achieves $1486\times$ geomean speedup over CPU and $11.87\times$ geomean speedup over the state-of-the-art at iso-area. Together, these results demonstrate compelling performance while scaling to large problem sizes (upwards of $2^{30}$ constraints) and maintaining small proof sizes (454-5 KB).
@@ I Introduction
Zero-Knowledge Proofs (ZKPs) are cryptographic protocols that allow a prover to convince one or more verifiers that a claim is correct without revealing any information beyond the validity of the claim. For example, a prover can demonstrate possession of a valid authentication key without disclosing the key itself. ZKPs have many compelling applications, including privacy-preserving authentication, trustworthy AI, blockchain scalability, and recently, private age verification [57]. A major limitation of state-of-the-art ZKP protocols is their high computational cost–generating a single proof on a CPU can take minutes to hours. As a result, there is growing interest in accelerating ZKPs, both with programmable solutions (e.g., GPUs [38, 23]), and ASICs [73, 10, 74, 54, 9].
@@
ZKP protocols continue to evolve and improve. Many have been proposed, offering different trade-offs in runtime, proof size, verifier time, setup (universal vs. application-specific), and security guarantees. A major family of ZKP protocols relies on the SumCheck protocol[37]. SumCheck-based protocols have $O(N)$ prover time, as compared to NTT-based protocols (a popular alternative[17, 14]), which require $O(N\log(N))$ prover complexity. The SumCheck protocol is highly flexible and is at the core of many classic and modern ZKP protocols [16, 55, 60, 6, 67, 2, 70], which have relatively fast prover time. While SumChecks were part of previously accelerated systems [61, 54, 9], they studied specific instantiations of the SumCheck protocol rather than the protocol family more generally. Another major distinguishing factor between protocols is the type of polynomial commitment scheme[24] that is used. Some of these protocols, such as Spartan [55]+Orion [67], use more lightweight Merkle tree commitments at the cost of large proofs (many MBs), while HyperPlonk [6] uses a more computationally demanding commitment scheme that produces significantly smaller proofs (a few kBs).
@@
ZKP protocols typically encode computations as arithmetic circuits (ACs) – directed acyclic graphs where each node (or gate) enforces a local algebraic constraint, such as addition ($f=a+b$) or multiplication ($f=a\cdot b$) over a finite field. We generally denote the number of gates by $N$. These circuits are then arithmetized by translating gates into polynomial constraints, which are later verified using polynomial checks for correctness. The SumCheck protocol is used to reduce checking $N$ gates to checking a single, randomized gate formula. A key feature of newer SumCheck-based protocols is the use of custom, high-degree gates that encapsulate more computation per gate by using higher-degree polynomial constraints (e.g., $f=a\cdot b^{5}+c^{2}$). This reduces the overall number of gates needed to represent a computation, resulting in large efficiency gains. The move to custom gates is exemplified in the Halo2 arithmetization language[69]. Halo2 enables a circuit designer to specify the types of gates they want to use per circuit.
@@
However, this level of customizability introduces significant hardware challenges. Unlike NTTs, which have fixed, regular dataflows, SumCheck computations vary with each gate’s structure and polynomial degree. High-degree polynomials require more modular multiplications and reductions, leading to increased bandwidth pressure, varied memory access patterns, and more complex scheduling. Designing efficient SumCheck hardware thus involves a trade-off: either build fixed-function units for specific gates, or develop a general, programmable architecture that can dynamically adapt to diverse polynomials without significantly compromising performance.
@@
zkSpeed [9] exemplifies the first approach, using dedicated hardware for Vanilla gates (additions and multiplications) within the HyperPlonk protocol. While this provides good performance for a fixed gate set, it limits flexibility: zkSpeed’s SumCheck unit cannot easily handle complex custom gates like those seen in the Halo2 arithmetization, reducing its applicability across general ZKP workloads. Furthermore, zkSpeed’s reliance on large on-chip scratchpads limits scalability, as its memory footprint grows with the number of gates, becoming impractical beyond $2^{24}$ gates, a scale increasingly common in emerging ZKP applications.
@@
Generalizable SumCheck cores are needed beyond ZKPs, too. In verifiable computing, SumCheck serves as the core verification mechanism in protocols that provide integrity guarantees without requiring zero knowledge. However, previous work in this area, including verifiable ASICs [61], has also relied on custom SumCheck units with fixed function, again restricting applicability and reuse.
@@
In this work, we adopt the second approach with zkPHIRE, introducing a programmable SumCheck accelerator that supports arbitrary polynomial structures and gate types. zkPHIRE addresses the irregularity and data reuse challenges intrinsic to SumCheck with fused compute pipelines, tree-based interconnects for efficient reductions, and flexible scheduling to handle diverse polynomial structures and degrees. This flexibility opens the door to a wide range of applications across ZKP and verifiable computing. We evaluate our programmable SumCheck unit on a diverse array of gate types, and then embed it into a full-system accelerator, zkPHIRE. zkPHIRE improves upon zkSpeed with more efficient memory organization and module abstraction that allows hardware to be reused across protocol steps. We conduct a protocol-level evaluation using the HyperPlonk with custom gates as a representative case study. Our key contributions are:
@@
  • The first hardware evaluation of high-degree gates that offer significant performance and scalability benefits.
@@ II-B Multi-Scalar Multiplication
Multi-Scalar Multiplication (MSM) is a key kernel used in polynomial commitments and relies on elliptic curve cryptography. MSMs involve computing a linear combination (i.e., dot product) of scalar and (2D or 3D) point pairs on an elliptic curve. Given scalars $k_{1},k_{2},\ldots,k_{n}$ and corresponding elliptic curve points $P_{1},P_{2},\ldots,P_{n}$, the goal is to efficiently compute $s=\sum_{i}k_{i}P_{i}$. MSMs are computationally intensive because each term $k_{i}P_{i}$ involves a point-scalar multiplication – a costly elliptic curve operation. MSMs constitute a significant portion of the total runtime in many zkSNARKs, including Groth16 and HyperPlonk. As a result, prior work has focused on accelerating MSMs using both application-specific and programmable platforms [10, 73, 36, 38, 23, 32, 74]. Rather than computing each scalar multiplication directly, optimized approaches such as Pippenger’s algorithm [49] are commonly employed. Pippenger’s algorithm improves efficiency by restructuring the computation to rely more heavily on faster operations like point addition and point doubling.
@@ II-C SumCheck
SumCheck [37] is a round-based interactive protocol where a prover $\mathcal{P}$ convinces a verifier $\mathcal{V}$ that the sum of a multivariate polynomial, $f(X_{1},\ldots,X_{\mu})$, over the boolean hypercube (all boolean values of $X_{i}\ldots X_{\mu}$) equals a claimed value $C$. The protocol proceeds in $\mu$ rounds, one per variable. In round $i$, $\mathcal{P}$ sends a univariate polynomial
@@
where $r_{1},\ldots,r_{i-1}$ are $\mathcal{V}$’s random challenges from prior rounds. $\mathcal{V}$ checks that $s_{i}(0)+s_{i}(1)$ equals the previous round’s claim (or the initial claim $C$ for round $1$). $\mathcal{V}$ then chooses a random field element $r_{i}$ and sends it to $\mathcal{P}$, who sets $X_{i}=r_{i}$ for the next round. After $\mu$ rounds, $\mathcal{V}$ evaluates $f$ at $(r_{1},\ldots,r_{\mu})$ and accepts if all checks pass.
ZKPs execute SumCheck over multilinear extensions (MLEs) – polynomials with degree at most $1$ in each variable. MLEs represent multivariate functions compactly and, in hardware, can be stored as flat lookup tables indexed by binary inputs. Throughout this paper, we use the terms “MLEs,” “MLE tables,” and “polynomials” interchangeably.
@@ II-C1 Polynomials and Gates
In ZKPs, polynomials encode and help enforce the correctness of the computation being proven. A popular encoding scheme is Plonk [14]. Plonk’s arithmetization scheme converts the underlying computation into an arithmetic circuit consisting of gates. Plonk circuits support additions, multiplications, conditionality, and equality checks. They can be represented by the composite polynomial
$f_{plonk}=q_{L}w_{1}+q_{R}w_{2}+q_{M}w_{1}w_{2}-q_{O}w_{3}+q_{C}$
The composite polynomial $f_{plonk}$ has 5 terms and a total of 8 unique, constituent polynomials. These constituent polynomials are the selectors (or enables) $q_{i}$ and witnesses $w_{i}$, and they are all multilinear. Notably, $f_{plonk}$ is not itself multilinear; rather, it is composed of multilinear polynomials.
Each of these polynomials encodes a function of a multivariate input vector $X=[X_{1},X_{2},\ldots,X_{\mu}]$, which encodes the gate index in binary. For example, if $\mu=3$, there are 8 gates, and gate index 3 would be represented by evaluating $f_{plonk}(1,1,0)$. For a multiply operation, we would set $q_{M}(1,1,0)=1$, $q_{O}(1,1,0)=1$ and the other enable polynomials to 0. Then, gate 3 would execute correctly if and only if $f_{plonk}(1,1,0)=0$, that is,
$w_{1}(1,1,0)\cdot w_{2}(1,1,0)-w_{3}(1,1,0)=0$
Representing circuits as polynomials lends itself naturally to SumCheck-based provers where we can check (with additional constraints) that each gate $f_{plonk}(X)$ evaluates to 0 and that
$\sum_{X}f_{plonk}(X)=0$
@@ II-C2 High-degree, Expressive Gates
The challenge of using Plonk gates is that all computations must be mapped to only what is supported by the Plonk polynomial, similar to mapping program code to a RISC ISA. In some cases, this results in circuits that require millions ($2^{20}$) of gates, and in turn, polynomials with millions of evaluations. Recent protocols and libraries [69, 6] support expressive, high-degree gates, which encode complex operations and significantly reduce gate count. For example, HyperPlonk’s Jellyfish gate captures common patterns such as hashing; across benchmarks, it reduces gate counts by up to 32× compared to less expressive Vanilla Plonk gates. Similarly, Halo2 uses custom gates for elliptic curve operations. However, these expressive gates increase the complexity of the SumCheck algorithm used to verify them.
@@ II-C3 SumCheck on Multilinear Polynomial Products
SumCheck over a single multilinear polynomial is relatively straightforward. Consider a polynomial $f(X_{1},X_{2},X_{3})$ with $\mu=3$ and its MLE table representation. In the first round, the prover computes the univariate polynomial $s_{1}(X_{1})$ by summing entries with $X_{1}=0$ and $X_{1}=1$, yielding two evaluations $s_{1}(0)$ and $s_{1}(1)$ sent to the verifier. The prover then computes an MLE Update by fixing $X_{1}$ to a random challenge $r_{1}$, typically generated by hashing the round evaluations (e.g., with SHA3). For each $(X_{2},X_{3})$, the entries $f(0,X_{2},X_{3})$ and $f(1,X_{2},X_{3})$ define a line in $X_{1}$, which is extrapolated to $r_{1}$: $f(r_{1},X_{2},X_{3})=f(0,X_{2},X_{3})\cdot(1-r_{1})+f(1,X_{2},X_{3})\cdot r_{1}$. This halves each MLE table before the next SumCheck round.
In recent SumCheck-based ZKP protocols [55, 67, 6, 3], we typically work with multi-term compositions of multilinear polynomials, like $f_{plonk}$ in Section II-C1. In these protocols, we are given only the constituent polynomials and their composition structure, and then must perform SumCheck over the composition. These SumChecks are naturally more complicated. In general, to run SumCheck on a $d$-degree polynomial, we require $d+1$ evaluations for $s_{i}(X_{i})$ after each $i^{th}$ round of SumCheck. For example, $f_{plonk}$’s overall degree is $3$, so $4$ evaluations are needed per SumCheck round. These evaluations must be computed for each individual MLE before computing term-wise products and summing across terms.
Figure 1 shows the dataflow for one polynomial product term. Each pair of entries in an MLE table actually represents the evaluations at $X_{i}=0$ and $1$; these values are then extended to $X_{i}=2,\ldots,d$, using the same formula as MLE Update (hence multilinear extension). Then, the extensions for all polynomials in a term are multiplied together. These products are summed down the MLE table, yielding $d+1$ values for a term. When there is more than one term, this process will be repeated for each term, with the $d+1$ evaluations summed across terms to form the final polynomial $s_{i}(X_{i})$ for round $i$.
We now present our generalizable SumCheck accelerator.